1F. Primes and the zeta function:

∏_{p} \frac{1}{1 − \frac{1}{p^{s}}} = ∑_{n=1}^{∞} \frac{1}{n^{s}}

7F. Newton’s Law of Gravitation:

F = \frac{Gm_{1}m_{2}}{r^{2}}

11F. Schrödinger’s Equation:

iℏ\frac{∂ψ}{∂t} = −\frac{ℏ^{2}}{2m}∇^{2}ψ + V ψ

12F?. Maxwell’s Equations in Vacuum

∇ · B = 0

∇ × B = \frac{1}{c} \frac{∂E}{∂t}

∇ · E = 0

∇ × E = −\frac{1}{c} \frac{∂B}{∂t}

13F. Einstein’s Equation:

E_{0} = mc^{2}

14F. Einstein’s General Relativity Equation:

R_{μν} − \frac{1}{2}Rg_{μν} + Λg_{μν} = 8πT_{μν}

16F. Heisenberg Uncertainty Principle:

ΔxΔp ≥ \frac{ℏ}{2}

19F?. Limiting ratio of Fibonacci numbers = golden mean = partial fraction expansion:

lim_{n→∞} \frac{F_{n+1}}{F_{n}} = \frac{1 + \sqrt{5}}{2} = 1 + \frac{1}{1 + \frac{1}{1+ \frac{1}{1+...}}}

lim_{n→∞} \frac{F_{n+1}}{F_{n}} = 1 + \frac{1}{1 + …}

68. Kepler’s laws: